# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 4.2.1. Many new examples of spaces $L_{\lambda 2^{n-1},n}$
    We first show that to have $\Omega=\Omega^{\prime}$ it is necessary and sufficient to find a matrix $M\in GL_{n}({\mathbb{F}}_{2})$ such that $\underline{V}M=\underline{V}^{\prime}$ and $\underline{U}M=\underline{U}^{\prime}$. In fact, by Proposition 4.1.i and 4.1.iii $\Omega=\Omega^{\prime}$ if, and only if, there exists $M\in GL_{n}({\mathbb{F}}_{2})$ such that $\underline{Q}^{\prime}=\underline{Q}M$. By uniqueness of 4.29, this is equivalent to have $(\underline{U}^{\prime})^{2}=(\underline{U}^{2})M$ and $(\underline{V}^{\prime})^{2}=(\underline{V}^{2})M$, and, since the entries of $M$ are elements of ${\mathbb{F}}_{2}$, it is equivalent to have that $\underline{U}^{\prime}=\underline{U}M$ and $\underline{V}^{\prime}=\underline{V}M$.
@@
    We have shown so far that applying the construction of Theorem 4.26 to the $n+1$-uple $\left(W_{1}(X+b^{2}),\dots,W_{n}(X+b^{2}),R+b\right)$ produces $(Q_{1}^{\prime},\dots,Q_{n}^{\prime})$ giving rise to a basis of $\Omega^{\prime}$ equivalent to $\Omega$. By applying part $(i)$ we conclude that the same $\Omega^{\prime}$ can only arise from $(W_{1}^{\prime},\dots,W_{n}^{\prime},R^{\prime})$ with
@@
    For every $\lambda\equiv 1\mod 2^{n}-2$ we can then construct infinite equivalence classes of spaces $L_{\lambda 2^{n-1},n}$ by picking $d=\frac{\lambda-1}{2^{n}-2}$ and all possible $n$-uples $(W_{1},\dots,W_{n})\in{\mathcal{W}}_{d}$.
@@
Even though in Theorem 4.26 we assumed $n\geq 3$, the construction of the $n$-uple $\underline{Q}$ in its proof makes sense also for $n=2$. Namely, for every pair $(W_{1},W_{2})\in{\mathcal{W}}_{d}$ one can set $V_{1}=W_{2}$, $V_{2}=W_{1}$ and fix a solution $\gamma$ to the congruences
@@
one produces a pair $(U_{1}^{2}+XV_{1}^{2},U_{2}^{2}+XV_{2}^{2})$ giving rise to a space $L_{2\lambda,2}$. We can show that this construction recovers all the spaces $L_{2\lambda,2}$. In fact, starting from pairs $\underline{U},\underline{V}$ satisfying Bezout’s identity $U_{1}V_{2}+V_{1}U_{2}=1$ (cf. the discussion following Corollary 4.22 to see that these define all possible spaces $L_{2\lambda,2}$), we have that
@@
We remark that, if $\Omega$ a space $L_{2^{n-1}\lambda,n}$ is arising from an element $(W_{1},\dots,W_{n},R)\in{\mathcal{W}}_{d}\times k[X]$ as in Theorem 4.26, then every étale pullback of $\Omega$ also arise from this construction. More precisely, let $S(X)\in k[X]$ with $S^{\prime}(X)=1$ and let $\sigma^{\star}(\Omega)$ be the pullback of $\Omega$ with respect to the morphism $\sigma\in\mathrm{End}({\mathbb{P}}^{1}_{k})$ induced by $X\mapsto S(X)$ (cf. Lemma 2.13). Then, we know by Proposition 4.17 that $(Q_{1}(S(X)),\dots,Q_{n}(S(X)))$ is the $n$-uple arising from Theorem 4.8 for $\sigma^{\star}(\Omega)$. One verifies that $Q_{i}(S(X))=(U_{i}(S(X))+RV_{i}(S(X))^{2}+XV_{i}(S(X))^{2}$, and as a result one proves that $\sigma^{\star}(\Omega)$ is the space arising from $(W_{1}\circ S,\dots,W_{n}\circ S,R+S-X)$.
@@ 5. Standard $L_{(p-1)p^{n-1},n}$ spaces and their subspaces
In this section, we analyze certain spaces $L_{\lambda p^{n-1},n}$ for $\lambda=p-1$ that appear implicitly in [6]. In that paper, they are instrumental to lift to characteristic zero certain local actions of elementary abelian groups and they have since been made explicit in [8] as well as in [10], in a generalized form. Due to their relevance, and for the fact that they were among the first examples of spaces $L_{\lambda p^{n-1},n}$ to be discovered, we choose to give these spaces a name and we call them standard $L_{(p-1)p^{n-1},n}$ spaces. In the first subsection, we explain how to construct these spaces using Theorem 4.8. Namely, we find sufficient conditions for a set of polynomial $Q_{1},\dots,Q_{n}$ as in the theorem to produce a standard space. In the second subsection, we show that we can construct all the subspaces of a given standard space using étale pullbacks of other standard spaces.
@@ 5.1. Standard spaces
A standard $L_{p^{n-1}(p-1),n}$ space $\Omega$ is a space $L_{p^{n-1}(p-1),n}$ such that there exists a $n$-uple $\underline{a}:=(a_{1},a_{2},\cdots,a_{n})\in k^{n}$ with $\Delta_{n}(\underline{a})\neq 0$ for which ${\mathcal{P}}(\Omega)=\langle a_{1},a_{2},\cdots,a_{n}\rangle_{{\mathbb{F}_{p}}}-\{0\}$.
@@
Then $\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ is a standard space by [10, Proposition 3.1]. This shows that, for every $\underline{a}\in k^{n}$ whose entries are ${\mathbb{F}_{p}}$-linearly independent, there exists a standard space whose set of poles is $\langle a_{1},a_{2},\cdots,a_{n}\rangle_{{\mathbb{F}_{p}}}-\{0\}$.
@@
with $\mu^{p^{n-1}}=\frac{-1}{\Delta_{n}(\underline{a})^{p-1}}$. Then we have $\omega_{i}=\frac{P_{i}}{P}dX$ where $P:=\Delta_{n}(\underline{Q})$ and $P_{i}:=\leavevmode\nobreak\ (-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$.
@@
Using [10, Proposition 3.1], we can write
@@
Each $\omega_{i}$ has $p(p-1)$ poles. The poles in ${\mathcal{P}}(\Omega)$ are all the elements of the multiplicative group ${\mathbb{F}}_{p^{2}}^{\times}$ and those in common between $\omega_{1}$ and $\omega_{2}$ are those elements of ${\mathbb{F}}_{p^{2}}$ that neither belong to the one-dimensional ${\mathbb{F}}_{p}$-vector space generated by $1$ nor to the one generated by $a_{2}$.
@@ 5.2. Subspaces of standard spaces via étale pullbacks
However, we have seen in Lemma 2.13 how to find spaces $L_{\lambda dp^{n},n}$ starting from spaces $L_{\lambda p^{n-1},n}$ via the pullback of the differential forms under suitable morphisms of degree $dp$. In this section, we characterize the subspaces of the standard spaces constructed in §5.1 as étale pullbacks of standard spaces of lower dimension. More precisely, let $n>1$, $1\leq t\leq n$ and $A:=\langle a_{1},\cdots,a_{n}\rangle_{{\mathbb{F}_{p}}}\subset k$ with $\underline{a}:=(a_{1},a_{2},\cdots,a_{n})$ such that $\Delta_{n}(\underline{a})\neq 0$. Let $A_{n-t}=\langle a_{1},...,a_{n-t}\rangle_{{\mathbb{F}}_{p}}$, $A_{n-t}^{s}=\langle a_{n-t+1},...,a_{n}\rangle_{{\mathbb{F}}_{p}}$ and let $P_{A_{n-t}}$ be the structural polynomial of $A_{n-t}$ (cf. Definition A.4). For every $n-t+1\leq i\leq n$, we set $\widetilde{a}_{i}:=P_{A_{n-t}}(a_{i})$. Then, by Lemma A.16 we have that $\Delta_{t}(\underline{\tilde{a}})\neq 0$ and hence $\widetilde{A_{n-t}^{s}}=\langle\tilde{a}_{n-t+1},...,...,\tilde{a}_{n}\rangle$ is a ${\mathbb{F}}_{p}$ vector space of dimension $t$. Let $\Omega(A)$ be the standard space associated with $A$ and let
@@
be the elements of its usual basis. Let $Q_{1},\dots,Q_{n}$ be the $n$-uple of polynomials arising from Theorem 4.8, which we know by proposition 5.3 to satisfy $Q_{i}=\mu\frac{\Delta_{2}(a_{i},X)}{X}$ with $\mu^{p^{n-1}}=\frac{-1}{\Delta_{n}(\underline{a})^{p-1}}$. Similarly, we let
be the elements of the usual basis of the standard $L_{p^{t-1}(p-1),t}$-space $\Omega(\widetilde{A_{n-t}^{s}})$. Finally, we let $\widetilde{Q_{n-t+1}},\dots,\widetilde{Q_{n}}$ be the $t$-uple of polynomials arising from Theorem 4.8, which satisfy $\widetilde{Q_{i}}=\tilde{\mu}\frac{\Delta_{2}(\tilde{a_{i}},X)}{X}$ with $\tilde{\mu}^{p^{t-1}}=\frac{-1}{\Delta_{t}(\tilde{a}_{n-t+1},\dots,\tilde{a}_{n})^{p-1}}$.
@@
[…]
@@
    The $t$-uple of polynomials arising from Theorem 4.8 for the basis $(\sigma^{\star}\tilde{\omega}_{n-t+1},...,\sigma^{\star}\tilde{\omega}_{n})$ of $\Omega^{s}=\sigma^{\star}(\Omega(\widetilde{A_{n-t}^{s}}))$ is given by
@@
    where $P_{{\mathcal{Q}}_{n-t}}$ is the structural polynomial of the ${\mathbb{F}}_{p}$-vector space ${\mathcal{Q}}_{n-t}:=\langle Q_{1},...,Q_{n-t}\rangle_{{\mathbb{F}}_{p}}$.
@@
    By Lemma A.5, we have that $P_{A_{n-t}}^{\prime}(X)=(-1)^{n-t}\Delta_{n-t}(a_{1},\dots,a_{n-t})^{p-1}\in k^{\times}$, so that the morphism $\sigma$ gives rise to an etale pullback as in Section 2.3. Since the degree of $P_{A_{n-t}}$ is $p^{n-t}$, it follows that $\sigma^{\star}(\Omega(\widetilde{A_{n-t}^{s}}))$ is a space $L_{(p-1)p^{n-1},t}$. For every $n-t+1\leq i\leq n$, we remark that $\tilde{\omega}_{i}=\frac{dF_{i}}{F_{i}}$ with
@@
    $=\prod_{(\epsilon_{1},\dots,\epsilon_{n})\in{{\mathbb{F}}_{p}^{n}}}(X-(a+\sum_{j=n-t+1}^{n}\epsilon_{j}a_{j}))^{\epsilon_{i}}.$
@@
    By Proposition 4.17, the $t$-uple of polynomials arising from Theorem 4.8 for the basis $(\sigma^{\star}\tilde{\omega}_{n-t+1},...,\sigma^{\star}\tilde{\omega}_{n})$ of $\sigma^{\star}(\Omega(\widetilde{A^{s}_{n-t}}))$ is $\eta\widetilde{Q_{n-t+1}}(P_{A_{n-t}}),\dots,\eta\widetilde{Q_{n}}(P_{A_{n-t}})$, with $\eta^{p^{t-1}}=\frac{1}{P_{A_{n-t}}^{\prime}}=\frac{(-1)^{n-t}}{\Delta_{n-t}(a_{1},\dots,a_{n-t})^{p-1}}$, as stated.
    From Proposition 4.6, we have that the $t$-uple of polynomials arising from Theorem 4.8 for the basis $\big{(}(-1)^{n-t}\omega_{n-t+1},...,(-1)^{n-t}\omega_{n}\big{)}$ of $\Omega^{s}$ is $\big{(}P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n})\big{)}$. We can apply Proposition 4.1.i to find that the $t$-uple of polynomials arising from Theorem 4.8 for the basis $\big{(}\omega_{n-t+1},...,\omega_{n}\big{)}$ of $\Omega^{s}$ is $\big{(}(-1)^{n-t}P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,(-1)^{n-t}P_{{\mathcal{Q}}_{n-t}}(Q_{n})\big{)}$. Since by $(i)$ we know that $\sigma^{\star}(\tilde{\omega}_{i})=\omega_{i}$, and Proposition 4.1.ii ensures the uniqueness of the $t$-uples of polynomials arising from Theorem 4.8 for the same basis, we have that
@@
Let $n=3$, and $t=2$. Choose $A^{\prime}:=\leavevmode\nobreak\ \langle\alpha,\beta\rangle_{{\mathbb{F}_{p}}}\leavevmode\nobreak\ \subset\leavevmode\nobreak\ k$ of dimension 2 and consider the corresponding standard $L_{p(p-1),2}$ space $\Omega(A^{\prime})$. Let $a_{1}\in k^{\times}$, and define $S(X):=X^{p}-a_{1}^{p-1}X$.
@@
Let $\Omega$ be a space $L_{\lambda p^{n-1},n}$ obtained as in Lemma 2.13 via an étale pullback of $\Omega_{1}$ a space $L_{\lambda_{1}p^{n-1},n}$. If $\Omega$ can also obtained via an étale pullback of $\Omega_{2}$, a $L_{\lambda_{2}p^{n-1},n}$ space with $\lambda_{1}<\lambda_{2}$ then $p|\lambda_{2}$.

### v2

@@ 4.2.1. Many new examples of spaces $L_{\lambda 2^{n-1},n}$
    We first show that to have $\Omega=\Omega^{\prime}$ it is necessary and sufficient to find a matrix $M\in GL_{n}({\mathbb{F}}_{2})$ such that $\underline{V}M=\underline{V}^{\prime}$ and $\underline{U}M=\underline{U}^{\prime}$. In fact, by Proposition 4.1.i and 4.1.iii $\Omega=\Omega^{\prime}$ if, and only if, there exists $M\in GL_{n}({\mathbb{F}}_{2})$ such that $\underline{Q}^{\prime}=\underline{Q}M$. By uniqueness of 4.28, this is equivalent to have $(\underline{U}^{\prime})^{2}=(\underline{U}^{2})M$ and $(\underline{V}^{\prime})^{2}=(\underline{V}^{2})M$, and, since the entries of $M$ are elements of ${\mathbb{F}}_{2}$, it is equivalent to have that $\underline{U}^{\prime}=\underline{U}M$ and $\underline{V}^{\prime}=\underline{V}M$.
@@
    We have shown so far that applying the construction of Theorem 4.25 to the $n+1$-tuple $\left(W_{1}(X+b^{2}),\dots,W_{n}(X+b^{2}),R+b\right)$ produces $(Q_{1}^{\prime},\dots,Q_{n}^{\prime})$ giving rise to a basis of $\Omega^{\prime}$ equivalent to $\Omega$. By applying part $(i)$ we conclude that the same $\Omega^{\prime}$ can only arise from $(W_{1}^{\prime},\dots,W_{n}^{\prime},R^{\prime})$ with
@@
    For every $\lambda\equiv 1\mod 2^{n}-2$ we can then construct infinite equivalence classes of spaces $L_{\lambda 2^{n-1},n}$ by picking $d=\frac{\lambda-1}{2^{n}-2}$ and all possible $n$-tuples $(W_{1},\dots,W_{n})\in{\mathcal{W}}_{d}$.
@@
Even though in Theorem 4.25 we assumed $n\geq 3$, the construction of the $n$-tuple $\underline{Q}$ in its proof makes sense also for $n=2$. Namely, for every pair $(W_{1},W_{2})\in{\mathcal{W}}_{d}$ one can set $V_{1}=W_{2}$, $V_{2}=W_{1}$ and fix a solution $\gamma$ to the congruences
@@
one produces a pair $(U_{1}^{2}+XV_{1}^{2},U_{2}^{2}+XV_{2}^{2})$ giving rise to a space $L_{2\lambda,2}$. We can show that this construction recovers all the spaces $L_{2\lambda,2}$. In fact, starting from pairs $\underline{U},\underline{V}$ satisfying Bezout’s identity $U_{1}V_{2}+V_{1}U_{2}=1$ (cf. the discussion following Proposition 4.20 to see that these define all possible spaces $L_{2\lambda,2}$), we have that
@@
We remark that, if $\Omega$ a space $L_{2^{n-1}\lambda,n}$ is arising from an element $(W_{1},\dots,W_{n},R)\in{\mathcal{W}}_{d}\times k[X]$ as in Theorem 4.25, then every étale pullback of $\Omega$ also arise from this construction. More precisely, let $S(X)\in k[X]$ with $S^{\prime}(X)=1$ and let $\sigma^{\star}(\Omega)$ be the pullback of $\Omega$ with respect to the morphism $\sigma\in\mathrm{End}({\mathbb{P}}^{1}_{k})$ induced by $X\mapsto S(X)$ (cf. Lemma 2.14). Then, we know by Proposition 4.19 that $(Q_{1}(S(X)),\dots,Q_{n}(S(X)))$ is the $n$-tuple arising from Theorem 4.10 for $\sigma^{\star}(\Omega)$. One verifies that $Q_{i}(S(X))=(U_{i}(S(X))+RV_{i}(S(X))^{2}+XV_{i}(S(X))^{2}$, and as a result one proves that $\sigma^{\star}(\Omega)$ is the space arising from $(W_{1}\circ S,\dots,W_{n}\circ S,R+S-X)$.
4.2.2. Non-existence of spaces $L_{p^{n-1},n}$ in odd characteristic
Another application of Proposition 4.20 is the following theorem, which settles the case $\lambda=1$ for every $n$ and every $p$.
Theorem 4.31.
Let $p>2$. Then, there exist no space $L_{p^{n-1},n}$.
Proof.
Let us prove this by contradiction. Assuming that such a space exist and applying Proposition 4.20 there exist a $n$-tuple $Q_{1},\dots,Q_{n}$ of polynomials of degree 1 satisfying
(4.32) $\left(\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})\cdot\Delta_{n}(\underline{Q})^{p-2}\right)^{(p-2)}=1.$
We have $Q_{i}=u_{i}X+v_{i}$ with $\Delta_{n}(\underline{u})\neq 0$ and so clearly $\underline{Q}^{\prime\prime}=0$, from which one sees that $\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})^{\prime}=0$ and that $\Delta_{n}(\underline{Q})^{\prime\prime}=0$. Equation (4.32) then becomes
$\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})\cdot\left(\Delta_{n}(\underline{Q})^{p-2}\right)^{(p-2)}=1.$
Using the fact that $\Delta_{n}(\underline{Q})^{\prime\prime}=0$, we can compute by induction on $j$ that
$\left(\Delta_{n}(\underline{Q})^{p-2}\right)^{(j)}=(-1)^{j}(j+1)!(\Delta_{n}(\underline{Q})^{\prime})^{j}\Delta_{n}(\underline{Q})^{p-2-j}\;\;\mbox{for every }\;\;1\leq j\leq p-2.$
In particular, we have
$\left(\Delta_{n}(\underline{Q})^{p-2}\right)^{(p-2)}=(\Delta_{n}(\underline{Q})^{\prime})^{p-2}$
and Equation (4.32) becomes
$\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})\cdot(\Delta_{n}(\underline{Q})^{\prime})^{p-2}=1.$
In particular, the polynomial $(\Delta_{n}(\underline{Q})^{\prime})^{p-2}$ is of degree zero. It follows that $\Delta_{n}(\underline{Q})^{\prime}$ is also of degree zero since we are assuming that $p\geq 3$. However, when we express this constant polynomial in terms of $\underline{u}$ and $\underline{v}$ we get
$\Delta_{n}(\underline{Q})^{\prime}=\det(\underline{u},\underline{u}^{p}X^{p}+\underline{v}^{p},\cdots,\underline{u}^{p^{n-1}}X^{p^{n-1}}+\underline{v}^{p^{n-1}})$
and by multilinearity of determinant, the term of higher degree of $\Delta_{n}(\underline{Q})^{\prime}$ is $\Delta_{n}(\underline{u})X^{p+p^{2}+\cdots+p^{n-1}}.$ Since $\Delta_{n}(\underline{u})\neq 0$ by hypothesis, we get a contradiction and the result follows. ∎
Remark 4.33.
When $n=2$, Theorem 4.31 corresponds to the first part of [Pagot02, Théorème 9] which says that, for $p>2$, there are no spaces $L_{p,2}$. This is enough to give an obstruction to lifts of certain $({\mathbb{Z}}/p{\mathbb{Z}})^{2}$ extensions, thanks to [GreenMatignon99, Theorem III.3.1] which implies that any lift of a $({\mathbb{Z}}/p{\mathbb{Z}})^{2}$–extension k[[z]]/k[[t]]k[\![z]\!]/k[\![t]\!] with a single ramification jump of conductor $m+1=p$ has equidistant branch locus. Following this strategy, one achieves [Pagot02, Théorème 13], which states that for $p>2$ lifts of $({\mathbb{Z}}/p{\mathbb{Z}})^{2}$–extensions with a single ramification jump of conductor $m+1=p$ do not exist.
Unfortunately, the generalization of [GreenMatignon99, Theorem III.3.1] to $m+1>p$ is not true, and one needs to consider also lifts with branch loci having more complicated geometries. Work in progress by Pagot addresses this issue by considering generalisations of spaces $L_{m+1,n}$ that correspond to these geometries, achieving a proof that lifts of $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$–extensions with a single ramification jump of conductor $m+1=p^{n-1}$ do not exist.
@@ 5. Standard $L_{(p-1)p^{n-1},n}$ spaces and their subspaces
In this section, we analyze certain spaces $L_{\lambda p^{n-1},n}$ for $\lambda=p-1$ that appear implicitly in [Matignon99]. In that paper, they are instrumental to lift to characteristic zero certain local actions of elementary abelian groups and they have since been made explicit in [Pagot02] as well as in [FresnelMatignon23], in a generalized form. Due to their relevance, and for the fact that they were among the first examples of spaces $L_{\lambda p^{n-1},n}$ to be discovered, we choose to give these spaces a name and we call them standard $L_{(p-1)p^{n-1},n}$ spaces. In the first subsection, we explain how to construct these spaces using Theorem 4.10. Namely, we find sufficient conditions for a set of polynomial $Q_{1},\dots,Q_{n}$ as in the theorem to produce a standard space. In the second subsection, we show that we can construct all the subspaces of a given standard space using étale pullbacks of other standard spaces.
@@ 5.1. Standard spaces
A standard $L_{p^{n-1}(p-1),n}$ space $\Omega$ is a space $L_{p^{n-1}(p-1),n}$ such that there exists a $n$-tuple $\underline{a}:=(a_{1},a_{2},\cdots,a_{n})\in k^{n}$ with $\Delta_{n}(\underline{a})\neq 0$ for which ${\mathcal{P}}(\Omega)=\langle a_{1},a_{2},\cdots,a_{n}\rangle_{{\mathbb{F}_{p}}}-\{0\}$.
@@
Then $\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ is a standard space by [FresnelMatignon23, Proposition 3.1]. This shows that, for every $\underline{a}\in k^{n}$ whose entries are $\mathbb{F}_{p}$}-linearly independent, there exists a standard space whose set of poles is $\langle a_{1},a_{2},\cdots,a_{n}\rangle_{{\mathbb{F}_{p}}}-\{0\}$.
@@
with $\mu^{p^{n-1}}=\frac{-1}{\Delta_{n}(\underline{a})^{p-1}}$. Then we have $\omega_{i}=\frac{P_{i}}{P}dX$ where $P:=\Delta_{n}(\underline{Q})$ and $P_{i}:=\penalty 10000\ (-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$.
@@
Using [FresnelMatignon23, Proposition 3.1], we can write
@@
Each $\omega_{i}$ has p(p1)p(p-1) poles. The poles in ${\mathcal{P}}(\Omega)$ are all the elements of the multiplicative group ${\mathbb{F}}_{p^{2}}^{\times}$ and those in common between $\omega_{1}$ and $\omega_{2}$ are those elements of ${\mathbb{F}}_{p^{2}}$ that neither belong to the one-dimensional ${\mathbb{F}}_{p}$-vector space generated by $1$ nor to the one generated by $a_{2}$.
@@ 5.2. Subspaces of standard spaces via étale pullbacks
However, we have seen in Lemma 2.14 how to find spaces $L_{\lambda dp^{n},n}$ starting from spaces $L_{\lambda p^{n-1},n}$ via the pullback of the differential forms under suitable morphisms of degree $dp$. In this section, we characterize the subspaces of the standard spaces constructed in §5.1 as étale pullbacks of standard spaces of lower dimension. More precisely, let $n>1$, $1\leq t\leq n$ and $A:=\langle a_{1},\cdots,a_{n}\rangle_{{\mathbb{F}_{p}}}\subset k$ with $\underline{a}:=(a_{1},a_{2},\cdots,a_{n})$ such that $\Delta_{n}(\underline{a})\neq 0$. Let $A_{n-t}=\langle a_{1},\dots,a_{n-t}\rangle_{{\mathbb{F}}_{p}}$, $A_{n-t}^{s}=\langle a_{n-t+1},\dots,a_{n}\rangle_{{\mathbb{F}}_{p}}$ and let $P_{A_{n-t}}$ be the structural polynomial of $A_{n-t}$ (cf. Definition A.4). For every $n-t+1\leq i\leq n$, we set $\widetilde{a}_{i}:=P_{A_{n-t}}(a_{i})$. Then, by Lemma A.17 we have that $\Delta_{t}(\underline{\tilde{a}})\neq 0$ and hence the ${\mathbb{F}}_{p}$-vector space $\widetilde{A_{n-t}^{s}}:=\langle\tilde{a}_{n-t+1},\dots,\tilde{a}_{n}\rangle_{{\mathbb{F}}_{p}}$ is $t$-dimensional. Let $\Omega(A)$ be the standard space associated with $A$ and let
$\omega_{i}=\sum_{(\epsilon_{1},\dots,\epsilon_{n})\in{{\mathbb{F}}_{p}^{n}}}\frac{\epsilon_{i}}{X-\sum_{j=1}^{n}\epsilon_{j}a_{j}}dX,\ 1\leq i\leq n$
be the elements of its usual basis. Let $Q_{1},\dots,Q_{n}$ be the $n$-tuple of polynomials arising from Theorem 4.10, which we know by proposition 5.3 to satisfy $Q_{i}=\mu\frac{\Delta_{2}(a_{i},X)}{X}$ with $\mu^{p^{n-1}}=\frac{-1}{\Delta_{n}(\underline{a})^{p-1}}$. Similarly, we let
@@
    We have that $).\Omega^{s}=\sigma^{\star}(\Omega(\widetilde{A_{n-t}^{s}}$)). More precisely, for every $n-t+1\leq i\leq n$, $\omega_{i}$ is equal to $\sigma^{\star}(\tilde{\omega}_{i})$, the pullback of $\tilde{\omega}_{i}$ with respect to the morphism ${\mathbb{P}}^{1}_{k}\overset{\sigma}{\rightarrow}{\mathbb{P}}^{1}_{k}$ induced by $X\mapsto P_{A_{n-t}}(X)$.
@@
    The $t$-tuple of polynomials arising from Theorem 4.10 for the basis $(\sigma^{\star}\tilde{\omega}_{n-t+1},\dots,\sigma^{\star}\tilde{\omega}_{n})$ of $)\Omega^{s}=\sigma^{\star}(\Omega(\widetilde{A_{n-t}^{s}}$)) is given by
@@
    where $P_{{\mathcal{Q}}_{n-t}}$ is the structural polynomial of the ${\mathbb{F}}_{p}$-vector space ${\mathcal{Q}}_{n-t}:=\langle Q_{1},\dots,Q_{n-t}\rangle_{{\mathbb{F}}_{p}}$.
@@
    By Lemma A.5, we have that $P_{A_{n-t}}^{\prime}(X)=(-1)^{n-t}\Delta_{n-t}(a_{1},\dots,a_{n-t})^{p-1}\in k^{\times}$, so that the morphism $\sigma$ gives rise to an etale pullback as in Section 2.3. Since the degree of $P_{A_{n-t}}$ is $p^{n-t}$, it follows that $)\sigma^{\star}(\Omega(\widetilde{A_{n-t}^{s}}$)) is a space $L_{(p-1)p^{n-1},t}$. For every $n-t+1\leq i\leq n$, we remark that $\tilde{\omega}_{i}=\frac{dF_{i}}{F_{i}}$ with
@@
    $=\prod_{(\epsilon_{1},\dots,\epsilon_{n})\in{\mathbb{F}}_{p}^{n}}(X-\sum_{j=1}^{n}\epsilon_{j}a_{j})^{\epsilon_{i}}.$
@@
    By Proposition 4.19, the $t$-tuple of polynomials arising from Theorem 4.10 for the basis $(\sigma^{\star}\tilde{\omega}_{n-t+1},\dots,\sigma^{\star}\tilde{\omega}_{n})$ of $)\sigma^{\star}(\Omega(\widetilde{A^{s}_{n-t}}$)) is $\eta\widetilde{Q_{n-t+1}}(P_{A_{n-t}}),\dots,\eta\widetilde{Q_{n}}(P_{A_{n-t}})$, with $\eta^{p^{t-1}}=\frac{1}{P_{A_{n-t}}^{\prime}}=\frac{(-1)^{n-t}}{\Delta_{n-t}(a_{1},\dots,a_{n-t})^{p-1}}$, as stated.
    From Proposition 4.6, we have that the $t$-tuple of polynomials arising from Theorem 4.10 for the basis $\big((-1)^{n-t}\omega_{n-t+1},\dots,(-1)^{n-t}\omega_{n}\big)$ of $\Omega^{s}$ is $\big(P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n})\big)$. We can apply Proposition 4.1.i to find that the $t$-tuple of polynomials arising from Theorem 4.10 for the basis $\big(\omega_{n-t+1},\dots,\omega_{n}\big)$ of $\Omega^{s}$ is $\big((-1)^{n-t}P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,(-1)^{n-t}P_{{\mathcal{Q}}_{n-t}}(Q_{n})\big)$. Since by $(i)$ we know that $\sigma^{\star}(\tilde{\omega}_{i})=\omega_{i}$, and Proposition 4.1.ii ensures the uniqueness of the $t$-tuples of polynomials arising from Theorem 4.10 for the same basis, we have that
@@
Let $n=3$, and $t=2$. Choose $A^{\prime}:=\penalty 10000\ \langle\alpha,\beta\rangle_{{\mathbb{F}_{p}}}\penalty 10000\ \subset\penalty 10000\ k$ of dimension 2 and consider the corresponding standard $L_{p(p-1),2}$ space $\Omega(A^{\prime})$. Let $a_{1}\in k^{\times}$, and define $S(X):=X^{p}-a_{1}^{p-1}X$.
@@
Let $\Omega$ be a space $L_{\lambda p^{n-1},n}$ obtained as in Lemma 2.14 via an étale pullback of $\Omega_{1}$ a space $L_{\lambda_{1}p^{n-1},n}$. If $\Omega$ can also be obtained via an étale pullback of $\Omega_{2}$, a $L_{\lambda_{2}p^{n-1},n}$ space with $\lambda_{1}<\lambda_{2}$ then $p|\lambda_{2}$.